OpenStudy (anonymous):

Use the distance formula to decide whether PQ=QR P(-4,3) Q(-2,1) R(0,-1)

OpenStudy (anonymous):

\[\sqrt{(y _2-y _1)^2+(x _2-x _1)}\]

OpenStudy (anonymous):

^2

OpenStudy (anonymous):

they are equal

OpenStudy (anonymous):

D\[d^2 = \sqrt{(x _{p} - x _{q})^2 + (y _{p}-y _{q})^2} for PQ = d^2 =\sqrt{-2^2 +2^2 } \ \[= \sqrt{8}\] so, now do the same thing for the segment RQ and compare...